a promisorry note will pay $60,000 at mayurity 20years from now. If you pay $20,000 for the note now, what rate compounded continously would you earn? Use the formula A=Pe^rt
Answer by stanbon(75887)

a promisorry note will pay $60,000 at maturity 20 years from now. If you pay $20,000 for the note now, what rate compounded continously would you earn? Use the formula A=Pe^rt
---------------------
60,000 = 20,000*e^(20r)
e^(20r) = 3
------------------
Take the natural log of both sides to get:
20r = ln3
r = (ln3)/20 = 0.05493
Rate = 5.493%
=====================
Cheers,
Stan H.